OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 √3i

OpenStudy (anonymous):

you need two numbers, \(r\) and \(\theta\)

OpenStudy (anonymous):

\(r=\sqrt{a^2+b^2}\) which in your case is \[r=\sqrt{3^2+(3\sqrt3)^2}\] what do you get?

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

how do i square 3 square root of 3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

square each one

OpenStudy (misty1212):

what is \(3^2\)?

OpenStudy (anonymous):

9

OpenStudy (misty1212):

and what is \(\sqrt3^2\)?

OpenStudy (anonymous):

3?

OpenStudy (misty1212):

yes so \[(3\sqrt3)^2=9\times 3=27\]

OpenStudy (misty1212):

now you have \[r=\sqrt{9+27}\] compute that (don't use decimals)

OpenStudy (anonymous):

square root of36, 6

OpenStudy (misty1212):

ok good so you have one of the two numbers you needed \(r=6\)

OpenStudy (misty1212):

now you need \(\theta\)

OpenStudy (misty1212):

you know how to find it?

OpenStudy (anonymous):

how do i get it

OpenStudy (anonymous):

nope

OpenStudy (misty1212):

hmm must be a trig class right?

OpenStudy (anonymous):

pre calc :/

OpenStudy (misty1212):

same thing

OpenStudy (misty1212):

use \[\cos(\theta)=\frac{a}{r}\\ \sin(\theta)=\frac{b}{r}\]

OpenStudy (misty1212):

in your case \[\cos(\theta)=\frac{-3}{6}=-\frac{1}{2}\\ \sin(\theta)=\frac{3\sqrt3}{6}=\frac{\sqrt3}{2}\]

OpenStudy (misty1212):

in other words, find the point on the unit circle \[(-\frac{1}{2},\frac{\sqrt3}{2})\] and that will give you \(\theta\) and you are done

OpenStudy (anonymous):

pi/3 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

okay one second

OpenStudy (misty1212):

it would be \(\frac{\pi}{3}\) except the first coordinate is \(-\frac{1}{2}\) not \(\frac{1}{2}\)

OpenStudy (anonymous):

so the cosine has to be negative but the sin is positive, second quadrant?

OpenStudy (misty1212):

yes

OpenStudy (anonymous):

2pi/3

OpenStudy (misty1212):

that is the one yes

OpenStudy (anonymous):

what do i do with r and theta

OpenStudy (misty1212):

put them in to \[r\left(\cos(\theta)+i\sin(\theta)\right)\]

OpenStudy (anonymous):

thank you!!!!!!!!!!!!!!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]